OpenStudy (anonymous):

Solve the following system. y = x + 3 4x + y = 18 (6, 3) (3, 6) (−3, 6) (3, −6)

OpenStudy (anonymous):

So this problem is actually pretty easy.... do you have any idea where to start?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Why dont you start by plugging in the pints? for example: im going to use (6, 3) (3) = (6) + 3 4(6) + (3) = 18 does that make any sense?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

btw thats not the answer just saying.....

OpenStudy (anonymous):

To solve this you can use substitution since this is a system you can substitute x+3 in for y Does that make sense?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so then you can simplfy 4x+ x + 3 = 18 And what do you get?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Good job!!! so now that you know what x is you can plus that in to either of the 2 equations to finish the problem and find y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So what did you get?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

OKay good... see it wasn't hard at all

OpenStudy (anonymous):

And this is random btw I love Snapback too..!!!

OpenStudy (anonymous):

thank you me too :)